# Formal Verification of Noisy Quantum Reinforcement Learning Policies

### v1

@@ Abstract.
Quantum reinforcement learning (QRL) aims to leverage quantum phenomena to develop sequential decision-making policies that achieve task objectives more effectively than their classical counterparts. Unlike reinforcement learning policies implemented on classical hardware, QRL policies are subject to additional uncertainty from quantum measurements and hardware imperfections. These uncertainties, including bit-flip, phase-flip, and depolarizing noise, can alter agent behavior and lead to unsafe policy behaviour. However, there do not exist approaches that provide systematic methods for verifying exactly whether trained QRL policies satisfy safety requirements under specific quantum noise conditions. We present QVerifier, a formal verification method that uses probabilistic model checking to rigorously analyze trained QRL policies, with and without modeled quantum noise, to verify satisfaction or violation with safety properties. Our method incrementally constructs a formal model of the policy–environment interaction by expanding all states reachable under the policy and assigning their transition probabilities using the model’s dynamics. Quantum measurement uncertainty and optional quantum hardware noise are incorporated directly into these probabilities. The resulting model is then verified using the Storm model checker to assess satisfaction of the safety properties. Experiments across multiple QRL environments show that our approach precisely quantifies how different models of quantum noise affect safety guarantees. The results reveal not only how specific noise models degrade policy performance, but also cases where certain noise can improve it. By enabling rigorous safety verification before deployment, QVerifier addresses a critical need: because access to quantum hardware is prohibitively expensive, pre-deployment verification is essential for any safety-critical use of quantum reinforcement learning. It targets a potential classical–quantum sweet spot: trained QRL policies that execute efficiently on quantum hardware, yet remain tractable for classical probabilistic model checking despite being too slow for real-time classical deployment.
@@ 1. Introduction
As quantum computing [nielsen2010quantum] transitions from theory to commercial reality [macquarrie2020emerging, 18166, 18276], it opens new opportunities for advancing machine learning [DBLP:journals/ncs/CerezoVHCC22, bano2025envisioning, 18277]. Quantum reinforcement learning (QRL) combines reinforcement learning (RL) with quantum computing to train sequential decision-making policies that may achieve task objectives more effectively than their classical counterparts by leveraging quantum phenomena such as superposition and entanglement [jerbi2021quantum, chen2024introduction, meyer2022survey, skolik2022quantum, DBLP:journals/corr/abs-2505-11862].
@@
In general, an RL agent aims to learn a near-optimal policy to achieve a fixed objective by taking actions and receiving feedback in the form of rewards and state observations from the environment [sutton2018reinforcement]. A policy is memoryless if it depends only on the current state, and it can be stochastic, assigning action probabilities from which the agent samples its next action.
@@
In QRL, policies are encoded by variational quantum circuits (VQCs) [skolik2022quantum], which are parameterized circuits optimized during training. A VQC takes a quantum encoding of the environment state and produces a quantum state, from which action probabilities are estimated via repeated measurements (shots). Each shot measures the circuit on the same encoded input, yielding a single classical outcome by projecting the quantum state onto one of its possible results; many such shots are needed to approximate the underlying action distribution. VQCs can be executed on quantum simulators or quantum hardware [wu2021application], though simulators become exponentially slower for large circuits [DBLP:journals/corr/abs-2410-12660].
@@
As research in QRL progresses into safety-critical systems [DBLP:conf/icassp/KimCPK25, niraula2021quantum, lv2025supply], it is crucial to ensure that these QRL policies are developed with safety in mind [bano2025envisioning, skolik2023robustness].
@@
However, trained policies may exhibit unsafe behavior [DBLP:conf/setta/GrossJJP22], such as collisions [DBLP:journals/access/BanL24, DBLP:journals/qmi/SinhaMK25], since rewards rarely capture complex safety requirements [DBLP:journals/aamas/VamplewSKRRRHHM22]. In QRL, additional uncertainty from stochastic measurements and device noise [DBLP:conf/hais/EscuderoAGB23, DBLP:journals/quantum/ArochKK24] on costly quantum computing hardware [DBLP:journals/sqj/MoguelRVBGM22] can further distort action selection.
@@
To resolve the safety issues mentioned above, formal verification methods like probabilistic model checking [DBLP:conf/setta/GrossJJP22] have been proposed to reason about the safety of RL on classical hardware [yuwangPCTL, DBLP:conf/formats/HasanbeigKA20, DBLP:conf/atva/BrazdilCCFKKPU14, DBLP:conf/tacas/HahnPSSTW19] and the correctness of quantum systems [gay2005probabilistic, baltazar2008quantum]. Model checking is not limited by properties that rewards can express but supports a broader range of properties that can be expressed by probabilistic computation tree logic (PCTL) [hansson1994logic], a temporal logic designed to express probabilistic branching-time properties. At its core, probabilistic model checking uses mathematical models to verify a system’s correctness with respect to a given safety property [baier2008principles].
@@
However, in the context of QRL policy verification, probabilistic model checking has not yet been applied, and there are limited tools for systematic comparisons of QRL policies [kaldari2025quantum, bowles2024better, kruse2025benchmarking].
@@
In this paper, we introduce a method, QVerifier, for model-checking trained QRL policies against safety properties while explicitly accounting for quantum uncertainty arising from measurements and quantum hardware noise. Given a trained QRL policy, a quantum noise model, and a formal environment model, we incrementally build a formal representation of the combined policy–noise-environment system, exploring only the states actually reachable under the trained QRL policy. We then use the probabilistic model checker Storm [DBLP:journals/sttt/HenselJKQV22] to verify whether the resulting formal model satisfies the desired safety property.
@@
Our main contributions are a method for formally verifying trained QRL policy behaviors in modeled environments, with and without quantum noise models, and an exact comparison approach for QRL and classical RL performance. By enabling rigorous safety verification before deployment, QVerifier addresses a critical need: access to quantum hardware is prohibitively expensive, making pre-deployment verification essential for safety-critical applications. QVerifier targets a potential sweet spot between classical and quantum computation, where trained QRL policies execute efficiently on quantum hardware but remain feasible to simulate classically for offline model checking. Moreover, the methodology underlying QVerifier could inform the design of future quantum model checkers.
@@ 2. Related Work
Research at the intersection of quantum computing, machine learning, and software engineering has grown rapidly in recent years [bano2025envisioning]. A key challenge across these areas is dealing with quantum oise, which affects both quantum circuits and the software built on top of them. Noise-aware quantum software engineering, therefore, forms an active research direction [16888].
@@
Probabilistic model checking has long been used to verify the correctness of quantum systems, such as quantum communication protocols [gay2005probabilistic] or quantum key distribution [baltazar2008quantum]. More recently, verification methods for quantum machine-learning classifier models have emerged. For example, Guan et al. propose a robustness verification method for quantum classifiers [DBLP:conf/cav/GuanFY21], and VeriQR provides automated robustness checking for such classifiers [DBLP:conf/fm/LinGFYS24]. However, none of these works apply probabilistic model checking to the verification of QRL policies.
@@
In RL on classical hardware, several works study the verification of trained RL policies [DBLP:conf/sigcomm/EliyahuKKS21, DBLP:conf/sigcomm/KazakBKS19, pmlr-v161-corsi21a, DBLP:journals/corr/DragerFK0U15, DBLP:conf/pldi/ZhuXMJ19, DBLP:conf/seke/JinWZ22]. A particularly relevant line of work, for this paper, uses probabilistic model checking to verify deep RL policies [DBLP:conf/setta/GrossJJP22, DBLP:conf/icaart/GrossS24, DBLP:journals/access/KwonK25]. COOL-MC [DBLP:conf/setta/GrossJJP22] introduced incremental model construction for RL policies and safety verification with the Storm model checker. Our work builds directly on this idea but extends it to handle QRL policies, including quantum measurements and hardware noise.
@@
Classical noise has also been studied in the context of classical RL, where perturbations typically affect the agent’s state observations rather than the policy’s inner workings [DBLP:conf/icaart/GrossS0023, DBLP:conf/aips/GrossS0023]. In contrast, QRL naturally introduces additional uncertainty from quantum-state measurements and noise arising from error-prone quantum circuit gates within the quantum circuit.
For improving robustness of QRL during training, RegQPG adds Lipschitz regularization to the policy gradient objective [DBLP:journals/corr/abs-2410-21117]. Unlike this training-time approach, we perform post-training verification to check whether a trained QRL policy satisfies a safety property under explicit hardware noise models (bit-flip, phase-flip, depolarizing, amplitude-damping).
Several works analyze how quantum noise affects QRL performance. Skolik et al. study realistic noise sources, including shot noise, coherent Gaussian noise, and decoherence, and show that their effects vary strongly with both the type and the magnitude of the noise [skolik2023robustness]. These insights motivate our goal of exactly verifying trained QRL policies under specific noise models, rather than studying robustness in aggregate.
@@
Some research modifies trained RL policy architecture by disrupting parts of it during verification [DBLP:conf/esann/GrossS24], but this differs from having QVC gate-level noise. Closely related to verification is testing, with extensive work on both testing quantum systems and using quantum computing for testing [18308, muqeet2024assessing, wang2025bqtmizer, oldfield2025faster, mendiluze2025quantum, wang2025quantum, araujo2025using].
@@ 3.2. Probabilistic Model Checking
Storm [DBLP:journals/sttt/HenselJKQV22] is a model checker. It enables the verification of properties in induced DTMCs, with reachability properties being among the most fundamental. These properties assess the probability of a system reaching a particular state. For example, one might ask, “Is the probability of the system reaching an unsafe state less than 0.1?” A property can be either satisfied or violated.
The general workflow for model checking with Storm is as follows (see also Figure 1): First, the system, in our setting, an induced DTMC, is modeled using a language such as PRISM [prism_manual]. Next, a property is formalized based on the system’s requirements. Using these inputs, the model checker Storm verifies whether the formalized property holds or fails within the model.
In probabilistic model checking, there is no universal “one-size-fits-all” solution [DBLP:journals/sttt/HenselJKQV22]. The most suitable tools and techniques depend on the specific input model and properties being analyzed. During model checking, Storm can proceed “on the fly”, exploring only the parts of the DTMC most relevant to the formal verification.
@@
Figure 1. General model checking workflow [DBLP:journals/sttt/HenselJKQV22]. First, the system needs to be formally modeled, for instance, via PRISM. Then, the requirements are formalized, for instance, via PCTL. Eventually, both are inputted into the model checker, like Storm, which verifies the property.
Properties verifiable via Storm include temporal logic formulas for DTMCs, defined on paths using PCTL [DBLP:journals/fac/HanssonJ94], a branching-time logic.
@@ 3.3.1. Entanglement
and entangled otherwise. Entangled states produce measurement correlations that are impossible to replicate with any classical system [rosset2013entangled].
@@ 3.3.2. Amplitude encoding
A common technique for encoding classical data vectors into quantum states is amplitude encoding [kaldari2025quantum]. Given a classical state $s\in S$ with $d$ components, amplitude encoding prepares a quantum state
@@ 3.4. Quantum REINFORCE
REINFORCE [williams1992simple] is a policy gradient RL algorithm that trains a parameterized policy $\pi_{\theta}$ by directly optimizing the expected cumulative reward (see Figure 2). Given the reward function $rew$ from Definition 1, the return $G_{t}$ from time step $t$ is the discounted sum of rewards along a trajectory
@@
In Quantum REINFORCE, the policy $\pi_{\theta}$ is implemented as a parameterized quantum circuit (see Figure 3), where $\theta$ represents quantum gate parameters [sequeira2023policy]. For state $s$, the circuit prepares a density matrix $\rho_{\theta}(s)$. Performing a computational basis measurement yields action probabilities
@@
The gradient $\nabla_{\theta}\log\pi_{\theta}(a\mid s)$ is computed using parameter-shift rules [mitarai2018quantum], and parameters are updated following standard REINFORCE.
@@ 4.1. Induced DTMC Construction
Given an MDP $M=(S,s_{0},Act,Tr,rew,AP,L)$ and a memoryless stochastic policy $\pi$, where $\pi(a\mid s)$ denotes the probability of selecting action $a$ in state $s$, we construct an induced DTMC $M^{\pi}$ that resolves all nondeterminism. The transition probability between any two states $s,s^{\prime}\in S$ is computed by marginalizing over all actions:
@@ 4.3. Limitations
Our approach supports verifying any memoryless stochastic quantum policy, but it has several constraints. We are restricted to discrete state and action spaces [DBLP:conf/setta/GrossJJP22]. The size of the parameterized quantum circuit also poses scalability limits [DBLP:journals/corr/abs-2410-12660]. Additionally, time-dependent noise effects such as thermal relaxation [jang2025balancing] cannot be modeled due to the Markovian property of our formalism [DBLP:conf/setta/GrossJJP22].
@@ 5.1. Setup
In the Freeway environment, the RL agent controls a chicken (up, down, or no operation) running across a highway filled with traffic to get to the other side. Every time the chicken gets across the highway, it earns a reward of one. An episode ends if the chicken gets hit by a car or reaches the other side. Each state is an image of the game’s state. Note that we use an abstraction of the original game, which sets the chicken into the middle column of the screen and contains fewer pixels than the original game, but uses the same reward function and actions [DBLP:conf/icaart/GrossS24].
@@ 5.1.2. Trained policies
In each environment, we train a classical REINFORCE agent and a quantum-enhanced REINFORCE agent with similar training conditions. Both agents are trained for the same number of epochs (always 10,000 epochs), and we use identical reward structures. While our focus is on these two types of REINFORCE agents, the approach generalizes to other memoryless QRL policies [meyer2022survey]. Note that we did not focus on archiving high-performing agents, as this paper focuses on the verification process itself. For details about the exact training source code, we refer to the technical setup.
@@ 5.2. Analysis
Finally, we observe an effect under low levels of amplitude-damping noise: in the Ski environment, performance increases and even surpasses the trained classical RL baseline by 27% (see Table 1 for comparison). This beneficial regime disappears as noise levels increase, suggesting that mild dissipation can occasionally stabilize or regularize quantum policies, whereas stronger noise ultimately degrades performance. These observations align with prior findings [jerbi2021quantum], which similarly report noise-induced performance enhancements.

### v2

@@ Abstract.
Quantum reinforcement learning (QRL) aims to leverage quantum phenomena to develop sequential decision-making policies that achieve task objectives more effectively than their classical counterparts. Unlike reinforcement learning policies implemented on classical hardware, QRL policies are subject to additional uncertainty from quantum measurements and hardware imperfections. These uncertainties, including bit-flip, phase-flip, and depolarizing noise, can alter agent behavior and lead to unsafe policy behaviour. However, there do not exist approaches that provide systematic methods for verifying exactly whether trained QRL policies satisfy safety requirements under specific quantum noise conditions. We present QVerifier, a formal verification method that uses probabilistic model checking to rigorously analyze trained QRL policies, with and without modeled quantum noise, to verify satisfaction or violation with safety properties. Our method incrementally constructs a formal model of the policy–environment interaction by expanding all states reachable under the policy and assigning their transition probabilities using the model’s dynamics. Quantum measurement uncertainty and optional quantum hardware noise are incorporated directly into these probabilities. When the policy was trained under matching noise conditions, this formal model is exact; when trained on physical hardware, it constitutes an idealized approximation, as unknown hardware noise prevents exact modeling. The resulting model is then verified using the Storm model checker to assess satisfaction of the safety properties. Experiments across multiple QRL environments show that our approach precisely quantifies how different models of quantum noise affect safety guarantees. The results reveal not only how specific noise models degrade policy performance, but also cases where certain noise can improve it. By enabling rigorous safety verification before deployment, QVerifier addresses a critical need: because access to quantum hardware is prohibitively expensive, pre-deployment verification is essential for any safety-critical use of quantum reinforcement learning. It targets a potential sweet spot between classical and quantum computation, where trained QRL policies could still be modeled classically for probabilistic model checking.
@@ 1. Introduction
As quantum computing [58] transitions from theory to commercial reality [48, 57, 1], it opens new opportunities for advancing machine learning [13, 10, 2]. Quantum reinforcement learning (QRL) combines reinforcement learning (RL) with quantum computing to train sequential decision-making policies that may achieve task objectives more effectively than their classical counterparts by leveraging quantum phenomena such as superposition and entanglement [38, 14, 52, 69, 15, 37].
@@
In general, an RL agent aims to learn a near-optimal policy to achieve a fixed objective by taking actions and receiving feedback in the form of rewards and state observations from the environment [71]. A policy is memoryless if it depends only on the current state, and it can be stochastic, assigning action probabilities from which the agent samples its next action.
@@
In QRL, policies are encoded by variational quantum circuits (VQCs) [69, 65, 67], which are parameterized circuits optimized during training. A VQC takes a quantum encoding of the environment state and produces a quantum state, from which action probabilities are estimated via repeated measurements (shots). Each shot measures the circuit on the same encoded input, yielding a single classical outcome by projecting the quantum state onto one of its possible results; many such shots are needed to approximate the underlying action distribution. VQCs can be executed on quantum simulators or quantum hardware [79], though simulating larger circuits becomes more and more difficult [16].
@@
As research in QRL progresses into safety-critical systems [42, 59, 47, 19], it is crucial to ensure that these QRL policies are developed with safety in mind [10, 70].
@@
However, trained policies may exhibit unsafe behavior [23], such as collisions [9, 68], since rewards does not capture complex safety requirements [73]. In QRL, additional uncertainty from stochastic measurements and device noise [72, 5] on costly quantum computing hardware [54] can further distort action selection.
@@
To resolve the safety issues mentioned above, formal verification methods like probabilistic model checking [23] have been proposed to reason about the safety of RL on classical hardware [77, 34, 12, 31] and the correctness of quantum systems [22, 8]. Model checking is not limited by properties that rewards can express but supports a broader range of properties that can be expressed by probabilistic computation tree logic (PCTL) [33], a temporal logic designed to express probabilistic branching-time properties. At its core, probabilistic model checking uses mathematical models to verify a system’s correctness with respect to a given safety property [7].
@@
However, in the context of QRL policy verification, probabilistic model checking has not yet been applied, and there are limited tools for systematic comparisons of QRL policies [40, 11, 43].
@@
In this paper, we introduce QVerifier, a method for model-checking trained QRL policies against safety properties while explicitly accounting for quantum uncertainty arising from measurements and modeled quantum hardware noise. Given a trained QRL policy, an assumed quantum noise model, and a formal environment model, we incrementally build a formal model of the combined policy–noise–environment system, exploring only the states reachable under the policy. When the policy was trained on a simulator, this model is exact; when trained on physical hardware, it constitutes an idealized approximation, as the true noise experienced during training and deployment may differ from the assumed model. After the building process, we use the probabilistic model checker Storm [35] to verify whether the resulting formal model satisfies the desired safety property.
@@
Our main contributions are a method for formally verifying trained QRL policy behaviors in modeled environments, with and without quantum noise models, and an exact comparison approach for QRL and classical RL performance. By enabling rigorous safety verification before deployment, QVerifier addresses a critical need: access to quantum hardware is prohibitively expensive, making pre-deployment verification essential for safety-critical applications. QVerifier targets a potential sweet spot between classical and quantum computation, where trained QRL policies could still be modeled classically for probabilistic model checking.
@@ 2. Related Work
Research at the intersection of quantum computing, machine learning, and software engineering has grown rapidly in recent years [10]. A key challenge across these areas is dealing with quantum noise, which affects both quantum circuits and the software built on top of them. Noise-aware quantum software engineering, therefore, forms an active research direction [3].
@@
Probabilistic model checking has long been used to verify the correctness of quantum systems, such as quantum communication protocols [22] or quantum key distribution [8]. More recently, verification methods for quantum machine-learning classifier models have emerged. For example, Guan et al. propose a robustness verification method for quantum classifiers [29], and VeriQR provides automated robustness checking for such classifiers [46]. However, none of these works apply probabilistic model checking to the verification of QRL policies.
@@
In RL on classical hardware, several works study the verification of trained RL policies [20, 41, 17, 18, 80, 39]. A particularly relevant line of work, for this paper, uses probabilistic model checking to verify deep RL policies [23, 26, 44]. COOL-MC [23] introduced incremental model construction for RL policies and safety verification with the Storm model checker. Our work builds directly on this idea but extends it to handle QRL policies, including quantum measurements and hardware noise.
@@
Classical noise has also been studied in the context of classical RL, where perturbations typically affect the agent’s state observations rather than the policy’s inner workings [25, 24]. In contrast, QRL naturally introduces additional uncertainty from quantum-state measurements and noise arising from error-prone quantum circuit gates within the quantum circuit.
For improving robustness of QRL during training, RegQPG adds Lipschitz regularization to the policy gradient objective [51]. Unlike this training-time approach, we perform post-training verification to check whether a trained QRL policy satisfies a safety property under explicit hardware noise models (bit-flip, phase-flip, depolarizing, amplitude-damping).
Several works analyze how quantum noise affects QRL performance. Skolik et al. study realistic noise sources, including shot noise, coherent Gaussian noise, and decoherence, and show that their effects vary strongly with both the type and the magnitude of the noise [70]. These insights motivate our goal of exactly verifying trained QRL policies under specific noise models, rather than studying robustness in aggregate.
@@
Some research modifies trained RL policy architecture by disrupting parts of it during verification [27], but this differs from having QVC gate-level noise. Closely related to verification is testing, with extensive work on both testing quantum systems and using quantum computing for testing [75, 56, 74, 60, 50, 76, 4].
@@ 3.2. Probabilistic Model Checking
Storm [35] is a model checker. It enables the verification of properties in induced DTMCs, with reachability properties being among the most fundamental. These properties assess the probability of a system reaching a particular state. For example, one might ask, “Is the probability of the system reaching an unsafe state less than 0.1?” A property can be either satisfied or violated.
The general workflow for model checking with Storm is as follows (see also Figure 1): First, the system, in our setting, an induced DTMC, is modeled using a language such as PRISM [62]. Next, a property is formalized based on the system’s requirements. Using these inputs, the model checker Storm verifies whether the formalized property holds or fails within the model.
In probabilistic model checking, there is no universal “one-size-fits-all” solution [35]. The most suitable tools and techniques depend on the specific input model and properties being analyzed. During model checking, Storm can proceed “on the fly”, exploring only the parts of the DTMC most relevant to the formal verification.
@@
Figure 1. General model checking workflow [35]. First, the system needs to be formally modeled, for instance, via PRISM. Then, the requirements are formalized, for instance, via PCTL. Eventually, both are inputted into the model checker, like Storm, which verifies the property.
Properties verifiable via Storm include temporal logic formulas for DTMCs, defined on paths using PCTL [32], a branching-time logic.
@@ 3.3.1. Entanglement
and entangled otherwise. Entangled states produce measurement correlations that are impossible to replicate with any classical system [64].
@@ 3.3.2. Amplitude encoding
A common technique for encoding classical data vectors into quantum states is amplitude encoding [40]. Given a classical state $s\in S$ with $d$ components, amplitude encoding prepares a quantum state
@@ 3.4. Quantum REINFORCE
REINFORCE [78] is a policy gradient RL algorithm that trains a parameterized policy $\pi_{\theta}$ by directly optimizing the expected cumulative reward (see Figure 2). Given the reward function $rew$ from Definition 1, the return $G_{t}$ from time step $t$ is the discounted sum of rewards along a trajectory
@@
In Quantum REINFORCE, the policy $\pi_{\theta}$ is implemented as a parameterized quantum circuit (see Figure 3), where $\theta$ represents quantum gate parameters [66]. For state $s$, the circuit prepares a density matrix $\rho_{\theta}(s)$. Performing a computational basis measurement yields action probabilities
@@
The gradient $\nabla_{\theta}\log\pi_{\theta}(a\mid s)$ is computed using parameter-shift rules [53], and parameters are updated following standard REINFORCE.
@@ 4.1. Induced DTMC Construction
Given an MDP $M$ and a memoryless stochastic policy $\pi$, where $\pi(a\mid s)$ denotes the probability of selecting action $a$ in state $s$, we construct an induced DTMC $M^{\pi}$ that resolves all nondeterminism. The transition probability between any two states $s,s^{\prime}\in S$ is computed by marginalizing over all actions
@@ 4.3. Limitations
Our approach supports verifying any memoryless stochastic quantum policy, but it has several constraints. We are restricted to discrete state and action spaces [23]. The size of the parameterized quantum circuit also poses scalability limits [16]. Additionally, time-dependent noise effects such as thermal relaxation [36] cannot be modeled due to the Markovian property of our formalism [23].
@@ 5.1. Setup
In the Freeway environment, the RL agent controls a chicken (up, down, or no operation) running across a highway filled with traffic to get to the other side. Every time the chicken gets across the highway, it earns a reward of one. An episode ends if the chicken gets hit by a car or reaches the other side. Each state is an image of the game’s state. Note that we use an abstraction of the original game, which sets the chicken into the middle column of the screen and contains fewer pixels than the original game, but uses the same reward function and actions [26].
@@ 5.1.2. Trained policies
In each environment, we train a classical REINFORCE agent and a quantum-enhanced REINFORCE agent with similar training conditions. Both agents are trained for the same number of epochs (always 10,000 epochs), and we use identical reward structures. While our focus is on these two types of REINFORCE agents, the approach generalizes to other memoryless QRL policies [52, 37]. Note that we did not focus on archiving high-performing agents, as this paper focuses on the verification process itself. For details about the exact training source code, we refer to the technical setup.
@@ 5.2. Analysis
Finally, we observe an effect under low levels of amplitude-damping noise: in the Ski environment, performance increases and even surpasses the trained classical RL baseline by 27% (see Table 1 for comparison). This beneficial regime disappears as noise levels increase, suggesting that mild dissipation can occasionally stabilize or regularize quantum policies, whereas stronger noise ultimately degrades performance. These observations align with prior findings [38], which similarly report noise-induced performance enhancements.